OpenStudy (anonymous):

differentiate y=ln2x/sqrtx using quotient rule

OpenStudy (phi):

\[ d \left( \frac{u}{v}\right) = \frac{ v\ du - u \ dv}{v^2} \]

OpenStudy (phi):

I assume you can at least work out what the denominator will be ?

OpenStudy (anonymous):

the question i have is differentiate with respect to x i used that formula you provided but they want the answer as the f(dx) qoutient rule

OpenStudy (tylerd):

that is quotient rule

OpenStudy (phi):

***i used that formula*** what did you get ?

OpenStudy (cggurumanjunath):